If all the energy of a 20-g bullet travelling at 1km/s is converted into heat, how many calories are generated?


1
Expert's answer
2022-05-02T13:42:08-0400
"Q=\frac{mv^2}{2}\\
=\frac{0.020*(1000)^2}{2}=10000\:\rm J=2390\: cal"